What is the pH of a 3% solution of atropine sulfate knowing that the molecular weight of atropine sulfate is 676.8 g/mol and the pKb of atropine base is 4.35.

Select one:

a. 2.9

b. 4.6

c. 1.9

d. 5.5

e. 11.1
Answer #1

(m/v) % of base = 3 %

mass of base = 3 g

moles of base = 3 / 676.8 = 4.43 x 10^-3 mol

volume = 100 mL = 0.1 L

concentration of base = 4.43 x 10^-3 / 0.1 = 0.0443 M

pOH = 1/2 (pKb - log C)

        = 1/2 (4.35 - log 0.0443)

       = 2.85

pH = 11.1